Grade 12th passGeneral Physics

A balloon has 5 g of air. A small hole is pierced into it. The air escapes at a uniform rate with a velocity of 4 cm/s. If the balloon shrinks completely in 2.5 seconds, Then the average force acting on the balloon is:
  1. 2 dyne
  2. 50 dyne
  3. 8 dyne
  4. 8 N
  5. 20 N

Yogiraj
10 Years agoGrade 12th pass

1 Answer

arun
ApprovedApproved Tutor Answer10 Years ago
in this just you have to apply newton’s second law of motion which states that force is equal to rate of change of momentum with respect to time therefore, 
using newton’s second law of motion
F = \frac{\mathrm{d} mv}{\mathrm{d} t}
as v is constant for this case that is v = 4 cm/s
\Rightarrow F = v\frac{\mathrm{d} m}{\mathrm{d} t}
and  \frac{\mathrm{d} m}{\mathrm{d} t} = \frac{5}{2.5}
\Rightarrow F = 4\times 2 = 8 dyne